# Limit Laws for Poincaré Recurrence and the Shrinking Target Problem

### v1

@@ 8. Proof of Variance Estimates
where $(a_{k})$ is any of the sequences $(M_{k})$, $(\hat{M}_{k})$, $(\mu(E_{k}))$ or $(\mu(\hat{E}_{k}))$. For the proofs of the variance estimates, the following lemma estimating the correlations of $(E_{n})$ and $(\hat{E}_{n})$ will be useful. The lemma is similar to [23, Lemma 4.2], [34, Lemma 3] and [35, Proposition 4.2], relaxing the multiple decorrelation conditions in those works by Condition D.3, but obtaining worse estimates as a result.
@@ 8.3. Proof of variance estimates for explicit radii
see for example [37, Theorem 1]. In particular, since $h\in L^{p}(m)$ for all $p\in(1,2]$, we have that
@@
for all $p\in(1,2]$. Since $F_{k}\to h$ $\mu$-a.e. and $(X,m)$ is a finite measure space, this implies that $\lim_{k\to\infty}\|F_{k}-h\|_{L^{1}(m)}=0$. By the Cauchy–Schwarz inequality, we obtain that $\lim_{k\to\infty}\|F_{k}-h\|_{L^{2}(\mu)}=0$.
@@
This implies (see e.g. [22, Proposition 4.12]) that $\hat{\sigma}_{n}^{-2}\hat{s}_{n}^{2}\to\mu(h)^{-1}h$ in $L^{1}$. In particular, we have that
@@ 9. Discussion of the Short Returns Estimate
We devote this section to a discussion of the short return estimates, (3.1) and (3.2), present in Sections 2.2 and 3. We compare (3.2) with known results in order to motivate (3.1). We assume that $\mu$ is absolutely continuous with respect to $m$. The method goes back to Collet [8] (see Corollary 2.4, Lemma 2.5 therein) who establishes an a.e. version of (3.1) for some non-uniformly hyperbolic maps on the interval. More precisely the estimate is of the following form: for $\mu$-a.e. $x$, there exist $C,N,p,\delta>0$ such that, for all $k>0$ and $l=1,\ldots,(\log k)^{1+p}$,
@@
Similar results can be found in [18, Lemma 4.1], [19, §6.2], [17, §13], which modify Collet’s argument.
@@
Suppose that $\mathop{}\!d\mu=h\mathop{}\!dm$ for a density $h$ such that $c\leq h\leq C$ for some constants $c,C>0$. Let $\hat{E}_{k,l}=\{x\in X:d(x,T^{l}x)<2r_{k}\}$ and suppose that there exists $\delta_{0}>0$ such that, for all $k,l>0$,
(9.2) $\mu(\hat{E}_{k,l})\lesssim\hat{M}_{k}^{\delta_{0}}.$
If $\hat{M}_{k}\lesssim(\log k)^{-\upsilon}$ for some $\upsilon>2\delta_{0}^{-1}$, then there exist $p_{0}>1$ and $\delta_{1},\eta_{1}>0$ such that, for all $k>0$ and $l=1,\ldots,(\log k)^{p_{0}}$,
(9.3) $\mu\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap T^{-l}B(x,r_{k})\bigr)>\hat{M}_{k}^{1+\delta_{1}}\bigr\}\lesssim\hat{M}_{k}^{\eta_{1}}.$
@@
Let $p_{0}\in(1,\delta_{0}\upsilon/2)$ and define
$\tilde{E}_{k}=\{x\in X:d(T^{l}x,x)\leq 2r_{k}\text{ for some }l=1,\ldots,(\log k)^{p_{0}}\}.$
Notice that $\tilde{E}_{k}\subset\bigcup_{l=1}^{(\log k)^{p_{0}}}\hat{E}_{k,l}$. Hence, by (9.2), $\mu(\tilde{E}_{k})\lesssim\hat{M}_{k}^{\delta}$ for some $\delta<\delta_{0}$, where we use that $\hat{M}_{k}\lesssim(\log k)^{-\upsilon}$ and $p_{0}<\delta_{0}\upsilon/2$. Let $\delta_{1}<\delta$ and define
$F_{k}:=\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap\tilde{E}_{k}\bigr)\geq\hat{M}_{k}^{1+\delta_{1}}\bigr\}.$
@@
Notice that if $x\in F_{k}$, then $\mathfrak{M}_{k}(x)\gtrsim m(B(x,r_{k}))^{-1}\hat{M}_{k}^{1+\delta_{1}}\geq C_{0}\hat{M}_{k}^{\delta_{1}}$ for some constant $C_{0}>0$ by the assumptions on the density $h$. A theorem of Hardy and Littlewood (see e.g. [37, Theorem 1]) states that
$m\{x:\mathfrak{M}_{k}(x)>c\}\lesssim\frac{\|\mathbf{1}_{E_{k}}h\|_{L^{1}(m)}}{c}.$
@@
$\mu(F_{k})\lesssim m\{x:\mathfrak{M}_{k}(x)\geq C_{0}\hat{M}_{k}^{\delta_{1}}\}\lesssim\mu(\tilde{E}_{k})\hat{M}_{k}^{-\delta_{1}}\lesssim\hat{M}_{k}^{\delta-\delta_{1}}.$
Let $\eta_{1}=\delta-\delta_{1}>0$. Observe that if $y\in B(x,r_{k})\cap T^{-l}B(x,r_{k})$ for some $k>0$ and $l=1,\ldots,(\log k)^{p_{0}}$, then $y\in T^{-l}B(y,2r_{k})$ and $y\in B(x,r_{k})\cap\tilde{E}_{k}$. Hence,
$\mu\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap T^{-l}B(x,r_{k})\bigr)>\hat{M}_{k}^{1+\delta_{1}}\bigr\}\leq\mu(F_{k})\lesssim\hat{M}_{k}^{\eta_{1}},$
@@
The estimate in (9.2) has been verified for several systems, see [18, Lemmas 3.4, 3.10] for Gibbs–Markov maps and non-uniformly expanding interval maps, and [23, Section 6.2] for Linear Anosov diffeomorphisms.
We now show that, by an application of Lemma 9.1, one can obtain (3.2), i.e.,
@@
for some $\delta>0$ and all $k>0$ and $l=1,\ldots,(\log k)^{p_{0}}$. Write $A_{k,l}(y)=\mu(B(y,r_{k})\cap T^{-l}B(y,r_{k}))$ and notice that $\mu(A_{k,l})\leq\mu(B(y,r_{k}))\lesssim\hat{M}_{k}$ by the assumptions on the density $h$. Then, for all $k>0$ and $l=1,\ldots,(\log k)^{p_{0}}$,

### v2

@@ 8. Proof of Variance Estimates
where $(a_{k})$ is any of the sequences $(M_{k})$, $(\hat{M}_{k})$, $(\mu(E_{k}))$ or $(\mu(\hat{E}_{k}))$. For the proofs of the variance estimates, the following lemma estimating the correlations of $(E_{n})$ and $(\hat{E}_{n})$ will be useful. The lemma is similar to [25, Lemma 4.2], [36, Lemma 3] and [38, Proposition 4.2], relaxing the multiple decorrelation conditions in those works by Condition D.3, but obtaining worse estimates as a result.
@@ 8.3. Proof of variance estimates for explicit radii
see for example [41, Theorem 1]. In particular, since $h\in L^{p}(m)$ for all $p\in(1,q]$ (where $q>2$), we have that
@@
for all $p\in(1,q]$. Since $F_{k}\to h$ $\mu$-a.e. and $(X,m)$ is a finite measure space, this implies that $\lim_{k\to\infty}\|F_{k}-h\|_{L^{2}(m)}=0$. By the Cauchy–Schwarz inequality, we obtain that $\lim_{k\to\infty}\|F_{k}-h\|_{L^{1}(\mu)}=0$.
@@
This implies (see e.g. [24, Proposition 4.12]) that $\hat{\sigma}_{n}^{-2}\hat{s}_{n}^{2}\to\mu(h)^{-1}h$ in $L^{1}$. In particular, we have that
@@ 9. Discussion of the Short Returns Estimate
We devote this section to a discussion of the short return estimates, (3.1) and (3.2), present in Sections 2.2 and 3.
9.1. Short returns for explicit radii
We first compare (3.2) with known results, in order to then motivate (3.1). We assume that $\mu$ is absolutely continuous with respect to $m$. The method goes back to Collet [8] (see Corollary 2.4, Lemma 2.5 therein) who establishes an a.e. version of (3.1) for some non-uniformly hyperbolic maps on the interval. More precisely the estimate is of the following form: for $\mu$-a.e. $x$, there exist $C,N,p,\delta>0$ such that, for all $k>0$ and $l=1,\ldots,(\log k)^{1+p}$,
@@ 9.1. Short returns for explicit radii
Similar results can be found in [20, Lemma 4.1], [21, §6.2], [19, §13], which modify Collet’s argument.
@@
Assume that there exists a constant $\delta_{0}>0$ such that, for all $\varepsilon>0$,
(9.2) $\mu\{x\in X:d(T^{l}x,x)<\varepsilon\}\lesssim\varepsilon^{\delta_{0}},$
and assume that $\mathop{}\!d\mu=h\mathop{}\!dm$ for a density $h\asymp 1$. Suppose that $r_{k}\lesssim(\log k)^{-\upsilon_{0}}$ for some $\upsilon_{0}>\delta_{0}^{-1}$. Then, there exist $\upsilon>1$ and $\delta,\eta>0$ such that, for all $k\geq 1$ and $l=1,\ldots,(\log k)^{\upsilon}$,
(9.3) $\mu\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap T^{-l}B(x,r_{k})\bigr)>\hat{M}_{k}^{1+\delta}\bigr\}\lesssim\hat{M}_{k}^{\eta}.$
@@
Fix $\upsilon\in(1,\delta_{0}\upsilon_{0})$. Let $\hat{E}_{k,l}=\{x:d(T^{l}x,x)<2r_{k}\}$ and notice that $\mu(\hat{E}_{k,l})\lesssim r_{k}^{\delta_{0}}\asymp\hat{M}_{k}^{\delta_{0}}$ by (9.2) and $h\asymp 1$. Now, let
$\tilde{E}_{k}=\{x\in X:d(T^{l}x,x)\leq 2r_{k}\text{ for some }j=1,\ldots,(\log k)^{\upsilon}\}$
and notice that $\tilde{E}_{k}\subset\bigcup_{j=1}^{(\log k)^{\upsilon}}\hat{E}_{k,l}$. Hence, $\mu(\tilde{E}_{k})\lesssim\hat{M}_{k}^{\delta_{1}}$ for some $0<\delta_{1}<\delta_{0}-\upsilon/\upsilon_{0}$, where we have used that $\hat{M}_{k}\asymp r_{k}\lesssim(\log k)^{-\upsilon_{0}}$ and $\upsilon<\delta_{0}\upsilon_{0}$. Let $\delta<\delta_{1}$ and define
$F_{k}:=\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap\tilde{E}_{k}\bigr)>\hat{M}_{k}^{1+\delta}\bigr\}.$
@@
Notice that if $x\in F_{k}$, then $\mathfrak{M}_{k}\gtrsim m(B(x,r_{k}))^{-1}\hat{M}_{k}^{1+\delta}>C_{0}\hat{M}_{k}^{\delta}$ for some constant $C_{0}>0$ since $m(B(x,r_{k}))\asymp r_{k}\asymp\hat{M}_{k}$. A theorem of Hardy and Littlewood (see e.g. [41, Theorem 1]) states that
$m\{x:\mathfrak{M}_{k}(x)>c\}\lesssim\frac{\|\mathbf{1}_{\tilde{E}_{k}}h\|_{L^{1}(m)}}{c}.$
@@
$\mu(F_{k})\lesssim m\{x:\mathfrak{M}_{k}(x)>C_{0}\hat{M}_{k}^{\delta}\}\lesssim\mu(\tilde{E}_{k})\hat{M}_{k}^{-\delta}\lesssim\hat{M}_{k}^{\delta_{1}-\delta}.$
Let $\eta=\delta_{1}-\delta>0$. Observe that if $y\in B(x,r_{k})\cap T^{-l}B(x,r_{k})$ for some $k\geq 1$ and $l=1,\ldots,(\log k)^{\upsilon}$, then $y\in T^{-l}B(y,2r_{k})$ and $y\in B(x,r_{k})\cap\tilde{E}_{k}$. Hence,
$\mu\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap T^{-l}B(x,r_{k})\bigr)>\hat{M}_{k}^{1+\delta}\bigr\}\leq\mu(F_{k})\lesssim\hat{M}_{k}^{\eta},$
@@
The estimate in (9.2) has been verified for several systems, see [20, Lemmas 3.4, 3.10] for Gibbs–Markov maps and non-uniformly expanding interval maps, and [25, Section 6.2] for Linear Anosov diffeomorphisms.
We now show that, by an application of Lemma 9.1, one can obtain (3.2) in Lemma 3.6, i.e.,
@@
for some $\delta>0$ and all $k>0$ and $l=1,\ldots,(\log k)^{\upsilon}$. Write $A_{k,l}(y)=\mu(B(y,r_{k})\cap T^{-l}B(y,r_{k}))$ and notice that $\mu(A_{k,l})\leq\mu(B(y,r_{k}))\lesssim\hat{M}_{k}$ by the assumptions on the density $h$. Then, for all $k>0$ and $l=1,\ldots,(\log k)^{\upsilon}$,
@@
9.2. Short returns for implicit radii
We restrict ourselves to the one-dimensional setting $X=[0,1]$. Suppose that $T\colon[0,1]\to[0,1]$ is a piecewise $C^{1}$ expanding map and that there exists a partition $\{C_{j}\}_{j=0}^{\infty}$ of $[0,1]$ (up to a $m$-null set) such that $T|_{C_{j}}\colon C_{j}\to[0,1]$ is a bijection. Define
$C_{j_{0},\ldots,j_{n}}=\bigcap_{i=0}^{n}T^{-i}C_{j_{i}}.$
Suppose further that $\mu$ is a Gibbs measure for an observable $\varphi$ (for existence see [4] for the finite branch setting and [40] for the countable branch setting). Then $\mu\ll\nu$ for a conformal measure $\nu$ with density $h$ satisfying $c^{-1}\leq h\leq c$ for some $c>0$. We can assume that the pressure of $\varphi$ is zero. Thus, there exists $D>0$ such that
$D^{-1}\leq\frac{\mu(C_{j_{0},\ldots,j_{n-1}})}{\exp(\varphi_{n})}\leq D,$
where $\varphi_{n}=\sum_{k=0}^{n-1}\varphi\circ T^{k}$. As before, we assume that $\mu$ satisfies Condition M.2 so that the radii functions $r_{k}\colon X\to(0,\infty)$ are well defined.
In order to obtain a short returns estimate for the implicit radii setting, we define the set $E_{k.l}$ differently than in the explicit setting. For a sequence $(M_{k})\subset(0,1]$, let $r_{k}^{L}(x)$ and $r_{k}^{R}(x)$ denote, respectively, quantities such that $L_{k}(x):=(x-r_{k}^{L}(x),x]$ and $R_{k}(x):=[x,x+r_{k}^{R}(x))$ both have measure $M_{k}$. It follows from the same argument establishing that $r_{k}\colon X\to(0,\infty)$ is Lipschitz continuous with Lipschitz constant $1$ that $r_{k}^{L},r_{k}^{R}$ are both Lipschitz continuous with $\lvert r_{k}^{L}\rvert_{\textup{Lip}}=\lvert r_{k}^{R}\rvert_{\textup{Lip}}=1$. Define
$E_{k,l}=\{x\in X:T^{l}x\in L_{k}(x)\cup R_{k}(x)\}.$
The following lemma adapts the argument used in the proof of [36, Lemma 6].
Lemma 9.2.
In addition to the above assumptions, suppose that $|T^{\prime}|\geq 3$. Then there exists $C>0$ such that, for all sequences $(M_{k})\subset(0,1]$ and all $k,l\geq 1$,
(9.4) $\mu(E_{k,l})\leq CM_{k}.$
Proof.
Fix $k,l\geq 1$ and $j_{0},\ldots,j_{l-1}$, and let $I=I_{j_{0},\ldots,j_{l-1}}:=E_{k,l}\cap C_{j_{0},\ldots,j_{l-1}}$. Let $p$ denote the unique periodic point $p\in I$ of period $l$. By the continuity of $T|_{C_{j_{0},\ldots,j_{l-1}}}$, we have that $I=(a,b)$ for some $a<p<b$. We will first establish that $\mu(T^{l}I)\lesssim M_{k}$. We consider two cases.
Case 1: $(T^{l})^{\prime}>0$ on $C_{j_{0},\ldots,j_{l-1}}$. Since $T$ is expanding with positive derivative, $T^{l}a<a<p<b<T^{l}b$ and $T^{l}I=(T^{l}a,T^{l}b)$. We focus on estimating $\mu((T^{l}a,p])$ since the corresponding argument for $\mu([p,T^{l}b))$ is similar. By the continuity of $r_{k}^{L}$, we have that $d(T^{l}a,a)=r_{k}^{L}(a)$ and
$d(T^{l}a,p)=d(T^{l}a,a)+d(a,p)=r_{k}^{L}(a)+d(a,p).$
At the same time, since $(T^{l})^{\prime}$ is continuous, there exists $a<\xi<p$ such that
$d(T^{l}a,p)=d(T^{l}a,T^{l}p)=(T^{l})^{\prime}(\xi)d(a,p).$
As a consequence, $((T^{l})^{\prime}(\xi)-1)d(a,p)=r_{k}^{L}(a)\leq r_{k}^{L}(p)+d(a,p)$, where we have used that $\lvert r_{k}^{L}\rvert_{\textup{Lip}}=1$. Therefore
$((T^{l})^{\prime}(\xi)-2)d(a,p)\leq r_{k}^{L}(p).$
Using that $(T^{l})^{\prime}\geq 3$, we have that $d(a,p)\leq r_{k}^{L}(p)$ and therefore (see Figure 1) that
$(T^{l}a,p]=(T^{l}a,a]\cup(a,p]\subset L_{k}(a)\cup L_{k}(p)$
and $\mu((T^{l}a,p])\leq 2M_{k}$.
[…]
Figure 1. The positions of $p,a$ and $T^{l}a$ when $(T^{l})^{\prime}>0$.
Case 2: $(T^{l})^{\prime}<0$ on $C_{j_{0},\ldots,j_{l-1}}$. In this case, $T^{l}I=(T^{l}b,T^{l}a)$, where $T^{l}a=a+r_{k}^{R}(a)$ and $T^{l}b=b-r_{k}^{L}(b)$. This implies (see Figure 2) that
$T^{l}I=(T^{l}b,T^{l}a)\subset(T^{l}b,b)\cup(a,T^{l}a)=L_{k}(b)\cup R_{k}(a)$
and $\mu(T^{l}I)\leq 2M_{k}$.
[…]
Figure 2. The interval $T^{l}I$ when $(T^{l})^{\prime}<0$.
We now prove that $\mu(E_{k,l})\lesssim M_{k}$. Using that $\nu$ is conformal and $\mu$ is a Gibbs measure, we have that
$\nu(T^{l}I_{j_{0},\ldots,j_{l-1}})=\int_{I_{j_{0},\ldots,j_{l-1}}}e^{-\varphi_{l}(x)}\mathop{}\!d\nu\gtrsim\mu(C_{j_{0}},\ldots,j_{l-1})^{-1}\nu(I_{j_{0},\ldots,j_{l-1}}).$
Consequently, using that $\mathop{}\!d\mu=h\mathop{}\!d\nu$ with $c^{-1}\leq h\leq c$, we have that
$\begin{split}\mu(I_{j_{0},\ldots,j_{l-1}})&\lesssim\nu(I_{j_{0},\ldots,j_{l-1}})\\ &\lesssim\nu(T^{l}I_{j_{0},\ldots,j_{l-1}})\mu(C_{j_{0},\ldots,j_{l-1}})\\ &\lesssim\mu(T^{l}I_{j_{0},\ldots,j_{l-1}})\mu(C_{j_{0},\ldots,j_{l-1}})\\ &\lesssim M_{k}\mu(C_{j_{0},\ldots,j_{l-1}}).\end{split}$
We conclude the proof by summing over $j_{0},\ldots,j_{l-1}$. ∎
Remark 11.
An analysis of the proof of Lemma 9.2 makes it clear that one can relax the conditions at the cost of a more complicated formulation:
  1. (1)
    Firstly, the conclusion holds for all $l\geq 1$ such that $|(T^{l})^{\prime}|\geq 3$. Furthermore, one can relax the condition that $|(T^{l})^{\prime}(x)|\geq 3$ to $(T^{l})^{\prime}(x)\geq 3$ for all $x$ such that $(T^{l})^{\prime}(x)>0$, with no condition on the magnitude of the derivative whenever it is negative.
  2. (2)
    Secondly, assume that there exist $C_{1},C_{2},\delta>0$ such that, for all $x\in X$, there exist $s_{2}(x)\leq s_{1}(x)$ such that $s_{2}(x)/s_{1}(x)\geq\delta$ and, for all $r>0$,
    $C_{1}r^{s_{1}(x)}\leq\mu(B(x,r))\leq C_{2}r^{s_{2}(x)}.$
    Then one can replace the condition $|(T^{l})^{\prime}|\geq 3$ in Lemma 9.2 with $|(T^{l})^{\prime}|\geq 2$ and instead obtain that
    $\mu(E_{k,l})\lesssim M_{k}^{\delta}.$
As a result, the following systems satisfy (9.4): the doubling map $T_{2}x=2x\pmod{1}$ equipped with the Lebesgue measure; $n$-doubling maps $T_{n}x=nx\pmod{1}$ with integer $n\geq 3$ equipped with a Gibbs measure satisfying Condition M.1; and the Gauss map $G\colon[0,1]\to[0,1]$ defined by
$G(x)=\begin{cases}0&x=0,\\ \frac{1}{x}\pmod{1}&x\in(0,1]\end{cases}$
equipped with the Gauss measure $\mathop{}\!d\mu_{G}=(\log 2)^{-1}(1+x)^{-1}\mathop{}\!dm$.
Remark 12.
If the measure $\mu$ satisfies Condition M.1 (and therefore Condition M.2), then there exist $C_{0},s_{0}>0$ such that $M_{k}\leq C_{0}r_{k}^{L}(x)^{s_{0}}$ and $M_{k}\leq C_{0}r_{k}^{R}(x)^{s_{0}}$ for all $x\in[0,1]$ and $k\geq 1$. We can then apply this result to the explicit radii problem: For $\varepsilon_{k}>0$, let $M_{k}=C_{0}\varepsilon_{k}^{s_{0}}$. Then, for all $k,l\geq 1$,
(9.5) $\mu\{x:d(x,T^{l}x)<\varepsilon_{k}\}=\mu\{x:d(x,T^{l}x)\in L_{k}(x)\cup R_{k}(x)\}\lesssim M_{k}^{\delta}\lesssim\varepsilon_{k}^{s_{0}\delta}.$
This implies that there exist constants $C,\eta>0$ such that, for all $l\geq 1$
$\mu\{x:d(x,T^{l}x)<\varepsilon\}\leq C\varepsilon^{\eta}.$
(Compare with [20, Lemma 3.4], which establishes the above estimate for $\mu=m$.)
If, in addition, there exists $s_{1}>0$ such that $\mu(B(x,r))\gtrsim r^{s_{1}}$ for all $x$ and $r>0$, then $\mu(B(x,\varepsilon_{k}))\gtrsim\varepsilon_{k}^{s_{1}}$ and $\hat{M}_{k}\gtrsim\varepsilon_{k}^{s_{1}}$. Hence, we obtain from (9.5) that
$\mu\{x:d(x,T^{l}x)<\varepsilon_{k}\}\lesssim\hat{M}_{k}^{\delta_{0}}$
for $\delta_{0}=s_{0}\delta/s_{1}$. We have therefore verified (9.2).
Lemma 9.3.
Suppose that there exist $C,\delta_{0}>0$ such that, for all sequences $(M_{k})\subset(0,1]$ and all $k,l\geq 1$,
(9.6) $\mu(E_{k,l})\leq CM_{k}^{\delta_{0}}.$
Assume now that $(M_{k})$ is a sequence such that $M_{k}\lesssim(\log k)^{-\upsilon_{0}}$ for some $\upsilon_{0}>\delta_{0}^{-1}$. Then there exist $\upsilon>1$ and $\delta,\eta>0$ such that, for all $k\geq 1$ and $l=1,\ldots,(\log k)^{\upsilon}$,
$\mu\bigl\{x\in X:\mu\bigl(B(x,r_{k}(x))\cap T^{-l}B(x,r_{k}(x))\bigr)>M_{k}^{1+\delta}\bigr\}\lesssim M_{k}^{\eta}.$
Proof.
Fix $\upsilon\in(1,\delta_{0}\upsilon_{0})$, let
$\tilde{E}_{k}=\{x\in X:d(T^{l}x,x)\in L_{k}(x)\cup R_{k}(x)\text{ for some }j=1,\ldots,(\log k)^{\upsilon}\},$
and notice that $\tilde{E}_{k}\subset\bigcup_{j=1}^{(\log k)^{\upsilon}}E_{k,l}$. Hence, by (9.6), $\mu(\tilde{E}_{k})\lesssim M_{k}^{\delta_{1}}$ for some $0<\delta_{1}<\delta_{0}-\upsilon/\upsilon_{0}$, where we have used that $\delta_{0}-\upsilon/\upsilon_{0}>0$ by the choice of $\upsilon$. Let $\delta<\delta_{1}$ and define
$F_{k}:=\bigl\{x\in X:\mu\bigl(B(x,r_{k}(x))\cap\tilde{E}_{k}\bigr)>M_{k}^{1+\delta}\bigr\}.$
Define the maximal function $\mathfrak{M}_{k}$ for $\mathbf{1}_{\tilde{E}_{k}}$ with respect to the measure $\mu$ by
$\mathfrak{M}_{k}(x)=\sup_{r>0}\frac{1}{\mu(B(x,r))}\int_{B(x,r)}\mathbf{1}_{\tilde{E}_{k}}\mathop{}\!d\mu.$
Notice that if $x\in F_{k}$, then $\mathfrak{M}_{k}\gtrsim\mu(B(x,r_{k}(x)))^{-1}M_{k}^{1+\delta}>C_{0}M_{k}^{\delta}$ for some constant $C_{0}>0$ since $\mu(B(x,r_{k}(x)))=M_{k}$. By Fefferman [12], we have that
$\mu\{x:\mathfrak{M}_{k}(x)>c\}\lesssim\frac{\|\mathbf{1}_{E_{k}}\|_{L^{1}(\mu)}}{c}.$
As a consequence,
$\mu(F_{k})\leq\mu\{x:\mathfrak{M}_{k}(x)>C_{0}M_{k}^{\delta_{2}}\}\lesssim\mu(\tilde{E}_{k})M_{k}^{-\delta_{2}}\lesssim M_{k}^{\delta_{1}-\delta}.$
Let $\eta=\delta_{1}-\delta>0$. We will now show that $B(x,r_{k}(x))\cap T^{-l}B(x,r_{k}(x))$ is a subset of $B(x,r_{k}(x))\cap\tilde{E}_{k}$: Let $y\in B(x,r_{k}(x))\cap T^{-l}B(x,r_{k}(x))$. Then $y,T^{l}y\in B(x,r_{k}(x))$. The potentially problematic case is when $y$ and $T^{l}y$ appear on different sides of $x$, say $y\in(x-r_{k}(x),x]$ and $T^{l}y\in[x,x+r_{k}(x))$ (the other case is treated similarly). Since $B(x,r_{k}(x))$ and $R_{k}(y)=[y,y+r_{k}^{R}(y))$ have the same measure and $y\in B(x,r_{k}(x))$, it follows that $x,x+r_{k}(x))\subset R_{k}(y)$ (see Figure 3).
[…]
Figure 3. The intervals $B(x,r_{k}(x))$ and $R_{k}(y)$.
Hence, $T^{l}y\in R_{k}(y)$ and $y\in B(x,r_{k}(x))\cap\tilde{E}_{k}$. In conclusion,
$\mu\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap T^{-l}B(x,r_{k})\bigr)>M_{k}^{1+\delta}\bigr\}\leq\mu(F_{k})\lesssim M_{k}^{\eta},$
as required. ∎
Similar to how (3.2) in Lemma 3.6 is established, one can establish (3.1) in Lemma 3.4.
10. Final Remarks
10.1. Improvements
The first natural improvement to Theorems 2.1 and 2.2 is establishing an error term, that is, establishing a Berry–Esseen theorem. A Berry–Esseen theorem for the shrinking target problem would imply that of the recurrence problem since one can keep track of the error term in the proofs of Theorems 2.1 and 2.2. Moreover, one can ask for other limit laws for the recurrence problem such as a LIL or ASIP. The martingale approach used here to establish an ASIP for the shrinking target problem is not suitable for the recurrence problem since $E_{k}\notin T^{-k}\mathcal{B}$. Another question is whether Condition D.3 holds for Axiom A systems equipped with Gibbs measures other than the SRB measure.